# Cartesian Nodal Lines and Magnetic Kramers Weyl Nodes in Spin-Split Antiferromagnets

### v1

@@ Abstract
Figure 2: (a-b) Blue and red spheres denote MKWNs; purple and green spheres represent Weyl nodes at generic positions, with $\eta>\eta_{c}$ in (a) and $\eta<\eta_{c}$ in (b). Topological charges and cone tilting details of the Weyl nodes are shown in the middle inset. (c-d) Solid (dotted) orange lines refer to Fermi arcs on the left (right) $x$-normal surface, and the black arrows denote their spin polarizations. The solid green rings represent the projections of bulk Fermi surface, with the chemical potential fixed at $\mu=0.2$. $\lambda_{\rm M}=0.4$ and $0.8$ in (c) and (d), respectively. Common parameters are $t=t_{z}=0$, $\lambda_{\rm so}=1.1$, and $\eta=1$.
[…]
Apart from the symmetry-enforced MKWNs, the interplay between the SOC and the magnetic exchange field can generate additional Weyl nodes. For instance, let us first assume $\lambda_{\rm so}>\lambda_{\rm M}$. In this case, when the anisotropic parameter falls below the critical value $\eta_{c}=\lambda_{\rm so}/(2\lambda_{\rm M})$, two additional pairs of Weyl nodes emerge along the high-symmetry $k_{z}$ lines that traverse $(k_{x},k_{y})=(0,\pi)$ or $(\pi,0)$, as illustrated in Fig. 2(b). These topological charges along with theirs positions (shown by the subscript) are given by
$\mathcal{C}_{(0,\pi,-\pi+k_{0})}$ $=$ $\mathcal{C}_{(\pi,0,\pi-k_{0})}=1,$
$\mathcal{C}_{(0,\pi,-k_{0})}$ $=$ $\mathcal{C}_{(\pi,0,k_{0})}=-1,$ (7)
@@
where $k_{0}=\arcsin(\eta/\eta_{c})$. As $\eta$ increases, two Weyl nodes of opposite topological charges, initially located on the same high-symmetry $k_{z}$ line, move toward each other, meet, and eventually annihilate when $\eta$ exceeds the critical value. For the case where $\lambda_{\rm so}<\lambda_{\rm M}$, additional Weyl nodes emerge, leading to a more complex scenario. A detailed discussion of this case is provided in the SM [97]. It might be worth mentioning that the low-energy linearly dispersive Weyl cones associated with the MKWNs exhibit up-down symmetry. In contrast, the Weyl cones emerging at generic positions are tilted and typically belong to the type-II class [102] if the parameter $\lambda_{\rm so}$ is much smaller than the hopping parameter $t_{z}$.
@@
To further characterize these Weyl nodes, we turn to the Fermi arcs which depict the iso-energy contours of surface states. Since they originate and terminate at the projections of two Weyl nodes possessing opposite topological charges [103, 104, 105, 106], they serve as a distinctive hallmark indicating the presence of Weyl nodes [107, 108]. In this system, the $C_{4z}\mathcal{T}$ symmetry imparts unique features to both the Fermi arcs and their associated spin textures. In particular, when only MKWNs exist, each surface hosts two distinct Fermi arcs, each spanning across half of the surface Brillouin zone, as depicted in Fig.2(c). The spin textures of these arcs are fully polarized, and there exists a relationship governed by $C_{4z}\mathcal{T}$ symmetry between the spin textures on $x$-normal and $y$-normal surfaces. Specifically, the spin polarizations align along the $y$ direction on $x$-normal surfaces and along the $x$ direction on $y$-normal surfaces. When additional Weyl modes are introduced, the connections between the projections of the Weyl nodes via the Fermi arcs become more complex, yet the symmetry pattern of the spin textures remains unaltered, as illustrated in Fig.2(d).
Figure 3: (a) Two representative band degeneracy configurations giving rise to Hall plateaus. The Zeeman field is along the $z$ direction. (b) Evolution of the Hall conductivity tensor $\sigma_{xy}$ (in units of $e^{2}/h$) as a function of $B_{z}$. Parameters are $t=t_{z}=0$, $\mu=0$, and $\lambda_{\rm M}=\eta=1$.
@@
Anomalous Hall effect.—As aforementioned, although the Berry curvature is divergently large near the band degeneracy, the anomalous Hall effect arising from the Berry curvature is entirely canceled due to symmetry constraints. This picture is of course altered if a Zeeman field along the $z$-direction, represented by $B_{z}\sigma_{z}$, is present. The Zeeman field breaks the symmetries $\mathcal{M}_{xy}$, $\mathcal{M}_{\bar{x}y}$ and $C_{4z}\mathcal{T}$, and thereby allows the Hall conductivity tensor $\sigma_{xy}$ to be finite regardless of whether SOC is present or not [97]. When SOC is absent, i.e., $\lambda_{\rm so}=0$, we find that the Zeeman field deforms the CNLs into two nodal rings located at the $k_{z}=0$ and $\pi$ planes which are protected by the preserved mirror symmetry $\mathcal{M}_{z}$, as shown in the left subfigure of Fig.3(a). Notably, except for these two nodal planes, all other $k_{z}$ planes are gapped, and their Chern numbers are $\mathcal{C}(k_{z})=+1$ ($-1$) when $B_{z}<0$ ($B_{z}>0$), which collectively give rise to a three-dimensional quantum anomalous Hall effect if the Fermi surface corresponds to the two nodal rings [109].
@@
When $\lambda_{\rm so}$ is finite and $\lambda_{\rm so}<\lambda_{\rm M}$, we find that Hall plateaus can also be observed if $B_{z}>\lambda_{\rm so}$ and $\eta>\eta_{c}+|B_{z}|/2\lambda_{\rm M}$ are simultaneously fulfilled. Under these conditions, those Weyl nodes whose $k_{z}$-components of the positions depend on $B_{z}$ are annihilated with each other, leaving behind eight Weyl nodes whose $k_{z}$-components are independent of $B_{z}$. These eight Weyl nodes are organized into four pairs, located at momentum planes with $k_{z}=\{\pm k_{c}$, ±$\pm(\pi-k_{c})$}, where $k_{c}=\arcsin(\lambda_{\rm so}/\lambda_{\rm M})$, as illustrated in the right subfigure of Fig.3(a). These Weyl nodes act as boundaries, separating gapped $k_{z}$ planes with $\mathcal{C}=1$ from those with $\mathcal{C}=-1$. Consequently, they give rise to a Hall plateau at $\sigma_{xy}=(1-\frac{4}{\pi}\arcsin\frac{\lambda_{\rm so}}{\lambda_{\rm M}})e^{2}/h$ [110, 111], as shown in Fig.3(b). We emphasize that here the mechanism is different from the realization of three-dimensional quantum Hall effect based on Landau levels, which generally requires a strong magnetic field [112, 113, 114].
Quantized circular photogalvanic effect.—When Weyl nodes of opposite topological charges are separated in energy and the Weyl cones belong to type-I class, they can induce a quantized circular photogalvanic effect (CPGE) within a specific range of optical frequencies [115, 116, 117]. Specifically, when the momentum-space surface $\boldsymbol{S}$, formed by the momenta involved in the optical transition, encloses a Weyl node or multiple Weyl nodes with charge $\mathcal{C}_{\boldsymbol{q}}$, the trace of the CPGE tensor quantizes to the net topological charge of the Weyl nodes [115], i.e.,
[…]
where $\beta_{0}=\pi e^{3}/h^{2}$. In this system, when SOC is present, since inversion symmetry and all mirror symmetries are broken, the MKWNs of opposite topological charges are separated in energy, as illustrated in Fig.4(a). Additionally, the MKWNs are untilted, suggesting that this system is ideal for the observation of quantized CPGE.
Remarkably, we find that this system can support CPGEs with higher quantized values even though $\mathcal{C}_{\boldsymbol{q}}=\pm 1$ for individual Weyl nodes. In a Kramers Weyl semimetal, ${\rm Tr}\left[\beta(\omega)\right]$ typically quantizes to ±$\pm i\beta_{0}$, as time-reversal symmetry does not relate different Kramers Weyl nodes [99, 100]. In contrast, here the $C_{4z}\mathcal{T}$ symmetry ensures that the Weyl nodes near high-symmetry points $\rm X$ and $\rm Y$ ($\rm R$ and $\rm K$) emerge at the same energy, leading to a possible quantization at ±$\pm 2i\beta_{0}$. In the isotropic hopping limit where $t=t_{z}$, even quantization at ±$\pm 3i\beta_{0}$ can be achieved, as illustrated in Fig.4(b).
@@
Figure 4: (a) Band structures along high-symmetry lines of the Brillouin zone for varying strengths of anisotropy in the magnetic exchange field. (b) Trace of the CPGE tensor (in units of $i\beta_{0}$). Dashed and solid lines correspond to $\mu_{1}=-2t-t_{z}$ and $\mu_{2}=-2t+t_{z}$, respectively, with colors matching the corresponding band structures in (a). Common parameters are $t=t_{z}=0.9$, $\lambda_{\rm so}=1$, and $\lambda_{\rm M}=0.8$.
@@
Discussions and conclusions.—We unveil that CNLs and MKWNs are two distinctive forms of band degeneracies that can naturally emerge in spin-split antiferromagnets. These band degeneracies are characterized not only by divergent Berry curvature but also by topological boundary states exhibiting unconventional patterns of spin textures. Importantly, we predict intriguing phenomena stemming from these degeneracies, such as strong or even quantized anomalous Hall effects driven by weak Zeeman fields and CPGEs with higher-quantized values. Our findings hold broad relevance for spin-split antiferromagnets and suggest that these materials offer a rich platform for exploring unconventional topological phases and the associated phenomena.
@@
Acknowledgements.— Z.-Y.Z. and Z.Y. would like to thank Chang Liu, Song-Bo Zhang and Jin-Xin Hu for helpful discussion. Z.-Y.Z., D.Z. and Z.Y. are supported by the National Natural Science Foundation of China (Grant No. 12174455), and Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023B1515040023). Z. W. is supported by the National Natural Science Foundation of China (Grant No. 12474264), Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2404007), National Key R&D Program of China (Grant No. 2022YFA1404103).
@@
The supplemental material contains detailed derivation for the properties associated with the two distinctive forms of band degeneracies: Cartesian nodal lines (CNLs) and magnetic Kramers Weyl nodes (MKWNs). Three sections are in order: (I) Band degeneracies in the spin-split band structure; (II) Realization of the tight-binding Hamiltonian; (III) Anomalous Hall effects induced by a Zeeman field.
@@
$E_{\pm}(\boldsymbol{k})$ $=$ ±$\pm\sqrt{\sin^{2}k_{x}(\lambda_{\rm so}-\lambda_{\rm M}\sin k_{z})^{2}+\sin^{2}k_{y}(\lambda_{\rm so}+\lambda_{\rm M}\sin k_{z})^{2}+\left[\lambda_{\rm so}\sin k_{z}+\eta\lambda_{\rm M}(\cos k_{x}-\cos k_{y})\right]^{2}}$ (S2)
@@
where the superscript $c/v$ represents conduction/valence band. The term $\epsilon_{ijl}$ represents the antisymmetric Levi-Civita symbol, where $i$, $j$, and $l$ are indices belonging to the set $\{x,y,z\}$. Additionally, summation over repeated indices is implied. The vecvtor $\boldsymbol{d}(\boldsymbol{k})=\lambda_{\rm so}\boldsymbol{l}(\boldsymbol{k})+\lambda_{\rm M}\boldsymbol{m}(\boldsymbol{k})$ with $\boldsymbol{l}(\boldsymbol{k})=(\sin k_{x},\sin k_{y},\sin k_{z})$ and $\boldsymbol{m}(\boldsymbol{k})=(-\sin k_{x}\sin k_{z},\sin k_{y}\sin k_{z},\eta(\cos k_{x}-\cos k_{y}))$.
@@
It is straightforward to find that band degeneracies appear at these positions: (1) along the momentum lines satisfying $|k_{x}|=|k_{y}|$ within the $k_{z}=0/\pi$ planes, and (2) along the high-symmetry $k_{z}$ lines passing through $(k_{x},k_{y})=(0/\pi,0/\pi)$. These nodal lines intersect orthogonally at the four $C_{4z}\mathcal{T}$-invariant momenta within the Brillouin zone, whose explicit positions are at $(0,0,0)$, $(0,0,\pi)$, $(\pi,\pi,0)$ and $(\pi,\pi,\pi)$. At each intersection, the nodal-line structure is analogous to the Cartesian coordinate system, thereby we refer to these nodal lines as Cartesian nodal lines (CNLs). The three components of the Berry curvature for this case are given by
@@
In three dimensions, the antisymmetric Hall conductivity tensor also has three independent components, $\sigma_{xy}$, $\sigma_{yz}$ and $\sigma_{zx}$. When only considering the contribution from the Berry curvature, their relation with the Berry curvature is given by[90]
@@
The existence of topological boundary states in a nodal-line semimetal is generally characterized by a quantized $\pi$ Berry phase or winding number defined on lines traversing the Brillouin zone, provided that the considered line has inversion symmetry or chiral symmetry. These two symmetries also provide an intuitive understanding of the topological boundary states in this system. For instance, if we consider open boundary conditions in the principal $x$-direction, whether there exist topological boundary states can be determined by analyzing the reduced one-dimensional Hamiltonian,
@@
Here, $\lambda_{1}(k_{z})=\lambda_{\rm M}\sin k_{z}$, $\lambda_{2}(k_{y},k_{z})=\lambda_{\rm M}\sin k_{y}\sin k_{z}$, and $\eta(k_{y})=\eta\cos k_{y}$; We have omitted the term $\varepsilon_{0}(\boldsymbol{k})\sigma_{0}$ since it only affects the dispersion but has no impact on the existence of topological boundary states; Furthermore, both $k_{y}$ and $k_{z}$ have been treated as parameters since they are good quantum numbers when considering topological boundary states on the $x$-normal surfaces. It is evident that the one-dimensional Hamiltonian $\mathcal{H}(k_{x})$ has neither inversion symmetry nor chiral symmetry. This simple fact indicates that the Hamiltonian does not have topological boundary states on the $x$-normal surfaces. The absence of topological boundary states on the $x$-normal surfaces can also be understood by noting that, when the CNLs are projected along the $x$-direction, there always exist two nodal lines whose projections overlap (see the nodal-line structure shown in Fig.1 of the main text). However, away from the principal axis direction, the projection of certain nodal lines no longer overlaps with those of others, leading to the emergence of topological surface states. Below we consider open boundary conditions along the $(\bar{1}10)$ direction as an illustrative example. In this case, the projections of these nodal lines—specially $(\pi,\pi,k_{z})$, $(k,k,\pi)$ and $(k,k,0)$—do not overlap with those of any other nodal lines. To determine the topological surface states of this case, we perform a coordinate transformation to simplify the analysis. Specifically, we rotate the $k_{x}$-$k_{y}$ plane about the $(0,0,k_{z})$ axis by $\pi/4$. Introduce $k_{1}=\frac{k_{x}-k_{y}}{\sqrt{2}}$, $k_{2}=\frac{k_{x}+k_{y}}{\sqrt{2}}$. The Hamiltonian in the new coordinate system reads
@@
$=$ $-\lambda_{\rm M}\left\{\sqrt{2}\sin k_{z}\cos\frac{k_{1}}{\sqrt{2}}\sin\frac{k_{2}}{\sqrt{2}}\sigma_{-}+\sin\frac{k_{1}}{\sqrt{2}}\Lambda(k_{2},k_{z})\left[\cos\theta(k_{2},k_{z})\sigma_{+}+\sin\theta(k_{2},k_{z})\sigma_{z}\right]\right\}.$
@@
Above, we have defined $\sigma_{-}=\frac{\sigma_{x}-\sigma_{y}}{\sqrt{2}}$, $\sigma_{+}=\frac{\sigma_{x}+\sigma_{y}}{\sqrt{2}}$, $\theta(k_{2},k_{z})=\arg(\sqrt{2}\sin k_{z}\cos\frac{k_{2}}{\sqrt{2}}+i2\eta\sin\frac{k_{2}}{\sqrt{2}})$, and $\Lambda(k_{2},k_{z})=\sqrt{2\sin^{2}k_{z}\cos^{2}\frac{k_{2}}{2}+4\eta^{2}\sin^{2}\frac{k_{2}}{\sqrt{2}}}$. We note that this new set of Pauli matrices $\{\sigma_{-},\sigma_{+},\sigma_{z}\}$ satisfies the same algebra as the standard Pauli matrices $\{\sigma_{x},\sigma_{y},\sigma_{z}\}$. Specifically, they obey the anticommutation relation $\{\sigma_{i},\sigma_{j}\}=2\delta_{ij}$, and the commutation relation $\{\sigma_{i},\sigma_{j}\}=2i\epsilon_{ijl}\sigma_{l}$, where $i$, $j$ and $l\in\{-,+,z\}$. Here, the Levi-Civita symbol $\epsilon_{ijl}$ is defined such that $\epsilon_{-+z}=\epsilon_{+z-}=\epsilon_{z-+}=1$ and $\epsilon_{+-z}=\epsilon_{-z+}=\epsilon_{z+-}=-1$.

### v3

@@ Abstract
Figure 3: Magnetic Kramers Weyl nodes and their associated topological surface states. (a) Schematic of MKWNs in Brillouin zone when $\eta>\eta_{c}$. Blue and red spheres represent MKWNs with opposite charges, illustrated in the middle inset. (b) Fermi arcs on the left (right) $x$-normal surface is denoted by solid (dotted) purple lines, and the black arrows denote their spin polarizations. The solid green rings represent the projections of bulk states at fixed energy $E=0.2$. Parameters are $t=t_{z}=0$, $\lambda_{\rm M}=0.4$, $\lambda_{\rm so}=1.1$, and $\eta=1$.
@@
Beyond the symmetry-enforced MKWNs, the interplay between the SOC and the magnetic exchange field gives rise to additional Weyl nodes. Consider the case $\lambda_{\rm so}>\lambda_{\rm M}$. When the anisotropic parameter falls below the critical value $\eta_{c}=\lambda_{\rm so}/(2\lambda_{\rm M})$, two additional pairs of Weyl nodes emerge along the high-symmetry $k_{z}$ lines that traverse $(k_{x},k_{y})=(0,\pi)$ or $(\pi,0)$ (see Note S1 in the SI). These topological charges along with their positions (shown by the subscript) are given by $\mathcal{C}_{(0,\pi,-\pi+k_{0})}=\mathcal{C}_{(\pi,0,\pi-k_{0})}=1$ and $\mathcal{C}_{(0,\pi,-k_{0})}=\mathcal{C}_{(\pi,0,k_{0})}=-1$, where $k_{0}=\arcsin(\eta/\eta_{c})$. As $\eta$ increases, each pair of oppositely charged Weyl nodes on the same high-symmetry $k_{z}$ line, approaches, meets, and eventually annihilates when $\eta$ exceeds the critical value. In the regime $\lambda_{\rm so}<\lambda_{\rm M}$, a richer pattern of additional Weyl nodes emerges; details are provided in the Note S1 in the SI. It is worth noting that the low-energy linearly dispersive Weyl cones associated with the MKWNs exhibit up-down symmetry (type-I), while the Weyl cones emerging at generic positions tend to be tilted and typically belong to the type-II class [104] if the parameter $\lambda_{\rm so}$ is much smaller than the hopping parameter $t_{z}$.
@@
To further characterize these Weyl nodes, we turn to the Fermi arcs which depict the iso-energy contours of surface states, connecting the surface projections of Weyl nodes with opposite topological charges [105, 106, 107, 108]. They serve as a distinctive hallmark indicating the presence of Weyl nodes [109, 110]. In this system, the $C_{4z}\mathcal{T}$ symmetry imparts unique features to both the Fermi arcs and their associated spin textures. In particular, when only MKWNs exist, each surface hosts two distinct Fermi arcs, each spanning across half of the surface Brillouin zone, as depicted in Fig. 3(b). These arcs are fully spin polarized, and spin textures on $x$- and $y$-normal surfaces are related by $C_{4z}\mathcal{T}$ symmetry: the spin polarizations align along the $y$ ($x$) direction on $x$ ($y$)-normal surfaces. When additional Weyl modes are introduced, the connectivity of the Fermi arcs becomes more intricate, yet the spin texture pattern dictated by $C_{4z}\mathcal{T}$ symmetry remains unaltered (see Note S1 in SI).
Figure 4: Anomalous Hall effects activated by a $z$-direction Zeeman field. (a) Two representative band-degeneracy configurations giving rise to Hall plateaus under a Zeeman field applied along the $z$ direction. Left: $\lambda_{\rm so}=0$, where the $z$-direction Zeeman field deforms the CNLs into two nodal rings. Right: $0<\lambda_{\rm so}<B_{z}$, where four pairs of Weyl nodes (spheres in red and blue) remain whose $k_{z}$ components are fixed and independent of $B_{z}$. In both panels, the light red (light blue) regions mark the range of $k_{z}$ for which each $k_{z}$ slice is a Chern insulator with $\mathcal{C}=1$ ($\mathcal{C}=-1$). (b) Evolution of the Hall conductivity tensor $\sigma_{xy}$ (in units of $e^{2}/h$) as a function of $B_{z}$. Parameters are $t=t_{z}=0$, $\mu=0$, and $\lambda_{\rm M}=\eta=1$.
@@
Anomalous Hall effect.—As aforementioned, although the Berry curvature is divergently enhanced near band degeneracies, the anomalous Hall effect arising from the Berry curvature is entirely canceled due to symmetry constraints. This scenario changes once a Zeeman field is applied along the $z$-direction, represented by $B_{z}\sigma_{z}$. The Zeeman field breaks the symmetries $\mathcal{M}_{xy}$, $\mathcal{M}_{\bar{x}y}$ and $C_{4z}\mathcal{T}$, thereby allowing a large Hall conductivity $\sigma_{xy}$ regardless of whether SOC is present (see Note S3 in SI). In the absence of SOC ($\lambda_{\rm so}=0$), the Zeeman field deforms the CNLs into two nodal rings located at the $k_{z}=0$ and $\pi$ planes, protected by the remaining mirror symmetry $\mathcal{M}_{z}$, as shown in the left panel of Fig. 4(a). All other $k_{z}$ slices are gapped, and their Chern numbers are $\mathcal{C}(k_{z})=+1$ (1-1) for $B_{z}<0$ ($B_{z}>0$), leading to a three-dimensional quantum anomalous Hall effect when the Fermi surface coincides with the two nodal rings [111].
@@
When $\lambda_{\rm so}$ is finite and $\lambda_{\rm so}<\lambda_{\rm M}$, we find that Hall plateaus can also emerge provided the conditions $B_{z}>\lambda_{\rm so}$ and $\eta>\eta_{c}+|B_{z}|/2\lambda_{\rm M}$ are met. Under these parameter conditions, those Weyl nodes whose $k_{z}$-components of the positions vary with $B_{z}$ annihilate in pairs, leaving behind eight Weyl nodes whose $k_{z}$-components are fixed and independent of $B_{z}$. These residual nodes form four pairs, located at momentum planes $k_{z}=\{\pm k_{c}$, $\pm(\pi-k_{c})\}$ with $k_{c}=\arcsin(\lambda_{\rm so}/\lambda_{\rm M})$, as illustrated in the right panel of Fig. 4(a). They delineate gapped $k_{z}$ planes with $\mathcal{C}=1$ from those with $\mathcal{C}=-1$, producing a Hall plateau with quantized value $\sigma_{xy}=(1-\frac{4}{\pi}\arcsin\frac{\lambda_{\rm so}}{\lambda_{\rm M}})e^{2}/h$ [112, 113], as depicted in Fig. 4(b). We emphasize that this mechanism is different from the conventional realization of three-dimensional quantum Hall effect based on Landau levels, which typically requires a strong magnetic field [114, 115, 116].
@@
Figure 5: Quantized circular photogalvanic effect. (a) Band structures along high-symmetry lines of the Brillouin zone for different values of anisotropy parameter $\eta$ in the magnetic exchange field. The two choices of Fermi energy are $\mu_{1}=-2t-t_{z}$ and $\mu_{2}=-2t+t_{z}$. (b) Trace of the CPGE tensor $\mathrm{Tr}[\beta(\omega)]$ (in units of $i\beta_{0}$). Dashed (solid) curves correspond to the Fermi energy set to $\mu_{1}$ ($\mu_{2}$), with colors matching the corresponding band structures (and $\eta$) in [(a)]. Common parameters are $t=t_{z}=0.9$, $\lambda_{\rm so}=1$, and $\lambda_{\rm M}=0.8$.
Quantized circular photogalvanic effect.—When Weyl nodes of opposite topological charges are energetically separated and form type-I Weyl cones, they can induce a quantized circular photogalvanic effect (CPGE) within a suitable frequency window [117, 118, 119]. Specifically, when the momentum-space surface $\boldsymbol{S}$, formed by the momenta involved in the optical transition, encloses Weyl nodes with net topological charge $\sum_{\boldsymbol{q}}\mathcal{C}_{\boldsymbol{q}}$, the trace of the CPGE tensor becomes quantized [117], i.e.,
${\rm Tr}\left[\beta(\omega)\right]=i\frac{e^{3}}{2h^{2}}\oint\boldsymbol{\Omega}\cdot d\boldsymbol{S}=i\beta_{0}\sum_{\boldsymbol{q}}\mathcal{C}_{\boldsymbol{q}},$ (4)
where $\beta_{0}=\pi e^{3}/h^{2}$. In our system, the presence of SOC breaks inversion and all mirror symmetries, resulting in MKWNs of opposite topological charge being separated in energy, as illustrated in Fig. 5(a). Additionally, the MKWNs are untilted, suggesting that this system is ideal for the observation of quantized CPGE.
Remarkably, our system can support CPGEs with enhanced quantization values even though $\mathcal{C}_{\boldsymbol{q}}=\pm 1$ for individual Weyl nodes. In Kramers Weyl semimetals, ${\rm Tr}\left[\beta(\omega)\right]$ typically quantizes to $\pm i\beta_{0}$, since time-reversal symmetry does not relate different Kramers Weyl nodes [88, 89]. In contrast, here the $C_{4z}\mathcal{T}$ symmetry ensures that the Weyl nodes near high-symmetry points—such as $\rm X$ and $\rm Y$, or $\rm R$ and $\rm K$—are energetically aligned, enabling higher quantization values such as $\pm 2i\beta_{0}$. In the isotropic limit where $t=t_{z}$, even larger quantization like $\pm 3i\beta_{0}$ can be achieved, as illustrated in Fig. 5(b).
@@
Discussions and conclusions.—We unveil that CNLs and MKWNs are two distinctive forms of band degeneracies that can naturally emerge in spin-split antiferromagnets. These band degeneracies are characterized not only by divergent Berry curvature but also by topological boundary states exhibiting unconventional patterns of spin textures. Importantly, they give rise to striking physical consequences, including strong—potentially quantized—anomalous Hall effects driven by weak Zeeman fields, as well as CPGEs with enhanced quantization.
In the absence of SOC, our model Hamiltonian exhibits band degeneracies consistent with those reported in MnTe2[23] upon performing a coordinate transformation, suggesting MnTe2 may serve as a potential candidate material for observing CNL-related physics (see Note S5 in SI). In addition, chiral antiferromagnets, such as Mn3IrGe and YMnO3 [103], may provide promising platforms for the realization of MKWNs. These distinctive forms of spin degeneracies can be directly probed by spin-resolved ARPES [23], and the associated large anomalous Hall response is expected to manifest as an enhanced Kerr rotation angle in magneto-optical Kerr effect (MOKE) measurements [120]. Finally, we note that because the formation of CNLs only requires three mutually orthogonal mirror planes, exact orthogonal CNLs naturally arise in cubic, tetragonal, and orthorhombic crystals. However, similar nodal degeneracies may also appear along high-symmetry lines in hexagonal and trigonal crystals, albeit in generally non-orthogonal configurations. Therefore, the mechanisms we identify are broadly applicable across five of the seven crystal systems. Overall, our findings highlight spin-split antiferromagnets as a rich platform for exploring unconventional topological phases and the associated phenomena.
@@
Acknowledgements.— Z.-Y.Z. and Z.Y. would like to thank Chang Liu, Song-Bo Zhang and Jin-Xin Hu for helpful discussion. We also want to thank the anonymous Referees for their insightful comments that help us improve the paper. This work was funded by the National Natural Science Foundation of China via grant No. 12174455, No. 12474264, Guangdong Basic and Applied Basic Research Foundation via grant No. 2023B1515040023, Guangdong Provincial Quantum Science Strategic Initiative via grant No. GDZX2404007, and National Key R&D Program of China via grant No. 2022YFA1404103.
@@
The supplemental material contains five sections: (I) Band degeneracies in the spin-split band structure; (II) Realization of the tight-binding Hamiltonian; (III) Anomalous Hall effects induced by a Zeeman field; (IV) Zeeman field effects on the local magnetic moments; (V) Connection to real-world materials.
@@
$E_{\pm}(\boldsymbol{k})$ $=$ $\pm\sqrt{\sin^{2}k_{x}(\lambda_{\rm so}-\lambda_{\rm M}\sin k_{z})^{2}+\sin^{2}k_{y}(\lambda_{\rm so}+\lambda_{\rm M}\sin k_{z})^{2}+\left[\lambda_{\rm so}\sin k_{z}+\eta\lambda_{\rm M}(\cos k_{x}-\cos k_{y})\right]^{2}}$ (S2)
@@
where the superscript $c/v$ represents conduction/valence band. The term $\epsilon_{ijl}$ represents the antisymmetric Levi-Civita symbol, where $i$, $j$, and $l$ are indices belonging to the set $\{x,y,z\}$. Additionally, summation over repeated indices is implied. The vector $\boldsymbol{d}(\boldsymbol{k})=\lambda_{\rm so}\boldsymbol{l}(\boldsymbol{k})+\lambda_{\rm M}\boldsymbol{m}(\boldsymbol{k})$ with $\boldsymbol{l}(\boldsymbol{k})=(\sin k_{x},\sin k_{y},\sin k_{z})$ and $\boldsymbol{m}(\boldsymbol{k})=(-\sin k_{x}\sin k_{z},\sin k_{y}\sin k_{z},\eta(\cos k_{x}-\cos k_{y}))$.
@@
It is straightforward to find that band degeneracies appear at these positions: (1) along the momentum lines satisfying $|k_{x}|=|k_{y}|$ within the $k_{z}=0/\pi$ planes, and (2) along the high-symmetry $k_{z}$ lines passing through $(k_{x},k_{y})=(0/\pi,0/\pi)$. These nodal lines intersect orthogonally at the four $C_{4z}\mathcal{T}$-invariant momenta within the Brillouin zone, whose explicit positions are at $(0,0,0)$, $(0,0,\pi)$, $(\pi,\pi,0)$, and $(\pi,\pi,\pi)$. At each intersection, the nodal-line structure is analogous to the Cartesian coordinate system, thereby we refer to these nodal lines as Cartesian nodal lines (CNLs). The three components of the Berry curvature for this case are given by
@@
In three dimensions, the antisymmetric Hall conductivity tensor also has three independent components: $\sigma_{xy}$, $\sigma_{yz}$ and $\sigma_{zx}$. When only considering the contribution from the Berry curvature, their relation with the Berry curvature is given by [94]
@@
The existence of topological boundary states in a nodal-line semimetal is generally characterized by a quantized $\pi$ Berry phase or winding number defined on lines traversing the Brillouin zone, provided that the considered line has inversion symmetry or chiral symmetry. These two symmetries also provide an intuitive understanding of the topological boundary states in this system. For instance, if we consider open boundary conditions in the principal $x$ direction, whether there exist topological boundary states can be determined by analyzing the reduced one-dimensional Hamiltonian,
@@
[…]
@@
$=$ $-\lambda_{\rm M}\left\{\sqrt{2}\sin k_{z}\cos\frac{k_{1}}{\sqrt{2}}\sin\frac{k_{2}}{\sqrt{2}}\sigma_{-}+\sin\frac{k_{1}}{\sqrt{2}}\Lambda(k_{2},k_{z})\left[\cos\theta(k_{2},k_{z})\sigma_{+}+\sin\theta(k_{2},k_{z})\sigma_{z}\right]\right\}$
$=$ $-\sqrt{2}\lambda_{\rm M}\sin k_{z}\cos\frac{k_{1}}{\sqrt{2}}\sin\frac{k_{2}}{\sqrt{2}}\ \sigma_{-}-\lambda_{\rm M}\Lambda(k_{2},k_{z})\sin\frac{k_{1}}{\sqrt{2}}\sigma_{g}(k_{2},k_{z}),$ (S10)
@@
Above, we have defined $\sigma_{\pm}\equiv\frac{\sigma_{x}\pm\sigma_{y}}{\sqrt{2}}$, $\Lambda(k_{2},k_{z})=\sqrt{2\sin^{2}k_{z}\cos^{2}\frac{k_{2}}{\sqrt{2}}+4\eta^{2}\sin^{2}\frac{k_{2}}{\sqrt{2}}}$, and $\theta(k_{2},k_{z})=\arg(\sqrt{2}\sin k_{z}\cos\frac{k_{2}}{\sqrt{2}}+i2\eta\sin\frac{k_{2}}{\sqrt{2}})$. We note that this new set of Pauli matrices $\{\sigma_{-},\sigma_{+},\sigma_{z}\}$ satisfies the same algebra as the standard Pauli matrices $\{\sigma_{x},\sigma_{y},\sigma_{z}\}$. Specifically, they obey the anticommutation relation $\{\sigma_{i},\sigma_{j}\}=2\delta_{ij}$, and the commutation relation $\sigma_{i},\sigma_{j}]=2i\epsilon_{ijl}\sigma_{l}$, where $i$, $j$ and $l\in\{-,+,z\}$. Here, the Levi-Civita symbol $\epsilon_{ijl}$ is defined such that $\epsilon_{-+z}=\epsilon_{+z-}=\epsilon_{z-+}=1$ and $\epsilon_{+-z}=\epsilon_{-z+}=\epsilon_{z+-}=-1$. Furthermore, in Eq. (S10), we mathematically group $\cos\theta(k_{2},k_{z})\sigma_{+}+\sin\theta(k_{2},k_{z})\sigma_{z}$ as $\sigma_{g}(k_{2},k_{z})$. It is obvious that $\sigma_{g}(k_{2},k_{z})$ is a unitary matrix that always anticommutes with $\sigma_{-}$. Hence, there must be another unitary matrix $\mathcal{S}(k_{2},k_{z})=-\sin\theta(k_{2},k_{z})\sigma_{+}+\cos\theta(k_{2},k_{z})\sigma_{z}$ that anticommutes with both $\sigma_{g}(k_{2},k_{z})$ and $\sigma_{-}$, completing the Pauli-like algebra.